OpenStudy (anonymous):

What is the maximum volume in cubic inches of an open box to be made from a 10-inch by 20-inch piece of cardboard by cutting out squares of equal sides from the four corners and bending up the sides? Your work must include a statement of the function and its derivative.

OpenStudy (anonymous):

@Lyrae @AlexandervonHumboldt2 @CausticSyndicalist @mathmate @ShadowLegendX @SolomonZelman @TheSmartOne @sleepyjess

OpenStudy (mathmate):

Have you started calculus yet?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think the dimensions are V=(20-2x)(10-2x)x

OpenStudy (mathmate):

 Yes your expression is correct! V(x)=x(20-2x)(10-2x)

OpenStudy (mathmate):

You will need to find V'(x), equate to zero and solve for x. One of the solutions, say x1 will be minimum,and the other will (x2) be a maximum. To distinguish between the two, you calculate V"(x). V(x1)>0 => minimum, and V(x2)<0 => maximum.

OpenStudy (mathmate):

* V"(x1)>0 => minimum, V"(x2)<0 => maximum.

OpenStudy (anonymous):

\[f'(x)=4x(x^2-15x+50)\]

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

V(x)=x(20-2x)(10-2x) =\(4x^3-60x^2+200x\) Then you can take derivative

OpenStudy (anonymous):

\[12x^2-120x+200\]

OpenStudy (anonymous):

its the derivative so then? @mathmate

OpenStudy (mathmate):

Your V'(x)=12x2−120x+200 is correct. Now, you need to find x by equating V'(x)=0, i.e. solve the quadratic to find the maximum/minimum.

OpenStudy (anonymous):